# TopoTune: a framework for generalized combinatorial complex neural networks

### v2

@@ 3 Motivation
The position paper (Veličković, 2022) proposed that any function over a higher-order domain can be computed via message passing over a transformed graph, but without specifying how to design GNNs that reproduce CCNNs. Later, (Hajij et al., 2023) proposed that, given a combinatorial complex $\mathcal{C}$ and a collection of neighborhoods $\mathcal{N_{C}}$, a message-passing GNN that runs over the augmented Hasse graph $\mathcal{G}_{\mathcal{N_{C}}}$ is equivalent to a specific CCNN as in (3) running over $\mathcal{C}$ using
  • $\mathcal{N_{C}}$ as collection of neighborhoods:
  • same intra- and inter-aggregations, i.e., $\bigoplus$=$\bigotimes$ = ⨂;
  • and no rank- and neighborhood-dependent message functions, i.e., $\psi_{\mathcal{N},\textrm{rk}(\cdot)}=\psi\;\forall\mathcal{N}\in\mathcal{N_{C}}$.
Retaining expressivity, but not generality.
The works in (Jogl et al., 2022a; b) show that GNNs over the augmented Hasse graph $\mathcal{G}_{\mathcal{N_{C}}}$ are as expressive (using the WL criterion) as CCNNs over $\mathcal{C}$, i.e., they are equally effective in distinguishing non-isomorphic graphs. These results are appealing because they suggest that some specific CCNNs (as described in eq. 3) can be implemented with standard graph libraries without loss of expressivity. A next work from the same authors generalize these ideas to a wider class of non-standard message-passing GNNs (Jogl et al., 2024). As the authors state, message-passing GNNs over $\mathcal{G}_{\mathcal{N_{C}}}$ cannot model all CCNNs over $\mathcal{C}$ because they do not explicitly differentiate among neighborhoods and cells of different ranks. Therefore, GNNs over augmented Hasse graphs cannot be exhaustively used as surrogates of CCNNs, if not on the expressivity side. Moreover, these works focus on leveraging graph expansion to simulate existing CCNNs rather than to explore new architectures.
@@
Hypergraph neural networks have long relied on graph expansions (Telyatnikov et al., 2023), which has allowed the field to leverage advances in the graph domain and, by extension, a much wider breadth of models (Papillon et al., 2023). Most hypergraph models are expanded into graphs using the star (Zhou et al., 2006; Solé et al., 1996), the clique (Bolla, 1993; Rodríguez, 2002; Gibson et al., 2000), or the line expansion (Bandyopadhyay et al., 2020). As noted by Agarwal et al. (2006), many hypergraph learning algorithms in fact correspond to either the clique or star expansions with an appropriate weighting function. Further implementations are well captured in the survey (Antelmi et al., 2023).
The success story of hypergraph neural networks motivates further research on new graph-based expansions that generalize and subsume current CCNNs. These expansions could go beyond augmented Hasse graphs to, at the same time, encompass current CCNNs and exploit progress in the GNN field. Therefore, returning to our core goal of accelerating and democratizing TDL while preserving its theoretical properties, we propose a two-part approach: a novel graph-based methodology able to generate general architectures (Section 4), and a lightweight software framework to easily and widely implement it (Section 5).
@@ 4 Generalized Combinatorial Complex Neural Networks
We propose Generalized Combinatorial Complex Neural Networks (GCCNs), a novel broad class of TDL architecture. GCCNs overcome the limitations of previous graph-based TDL architectures by leveraging the notions of strictly augmented Hasse graphs and per-rank neighborhoods.
Figure 3: Ensemble of strictly augmented Hasse Graphs. Given a complex $\mathcal{C}$ with neighborhood structure including both incidence and upper adjacency (left), this graph projection (right) produces one augmented Hasse graph for each neighborhood.
@@
This graph expansion method (see Fig. 3) extends from the the established definition of an augmented Hasse graph (see Fig. 2). Specifically, given a combinatorial complex $\mathcal{C}$ and a collection of neighborhood functions $\mathcal{N_{C}}$, we expand it into $|\mathcal{N_{C}}|$ graphs, each of them representing a neighborhood $\mathcal{N}\in\mathcal{N}_{\mathcal{C}}$. In particular, the strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ of a neighborhood $\mathcal{N}\in\mathcal{N}_{\mathcal{C}}$ is a directed graph whose nodes $\mathcal{C}_{\mathcal{N}}$ and edges $\mathcal{E}_{\mathcal{N}}$ are given by:
@@
Following the same arguments from Hajij et al. (2023), a GNN over the strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}$ induced by $\mathcal{N}$ is equivalent to a CCNN running over $\mathcal{C}$ and using $\mathcal{N_{C}}=\{\mathcal{N}\}$ up to the (self-)update of the cells in $\mathcal{C}/\mathcal{C}_{\mathcal{N}}$.
@@
The standard definition of adjacencies and incidences given in Section 2 implies that they are applied to each cell regardless of its rank. For instance, consider a combinatorial complex of dimension two with nodes (0-cells), edges (1-cells), and faces (2-cells).
  • Employing the down incidence $\mathcal{N}_{I,\downarrow}$ as in (1) means the edges must exchange messages with their endpoint nodes, and faces must exchange messages with the edges on their sides. It is impossible for edges to exchange messages while faces do not.
  • Employing the up adjacency $\mathcal{N}_{A,\uparrow}$ as in (2) means the nodes must exchange messages with other edge-connected nodes, and edges must exchange messages the other edges bounding the same faces. It is impossible for nodes to exchange messages while edges do not.
This limitation increases the computational burden of standard CCNNs while not always increasing the learning performance, as we will show in the numerical results. For this reason, we introduce per-rank neighborhoods, depicted in Fig. 4. Formally, a per-rank neighborhood function $\mathcal{N}^{r}$ is a neighborhood function that, regardless of its definition, maps a cell $\sigma$ to the the empty set if $\sigma$ is not a $r$-cell (i.e., a cell of rank $r$). For example, the up/down $r$-incidences $\mathcal{N}^{r}_{I,\uparrow}$ and $\mathcal{N}^{r}_{I,\downarrow}$ are defined as
$\mathcal{N}^{r}_{I,\uparrow}(\sigma)=\left\{\begin{aligned} &\{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=\textrm{rk}(\sigma)+1,\sigma\subset\tau\}\textrm{ if rk}(\sigma)=r\\ &\emptyset\textrm{ otherwise}\end{aligned}\right.,$ (6)
$\mathcal{N}^{r}_{I,\downarrow}(\sigma)=\left\{\begin{aligned} &\{\tau\in\mathcal{C}\,|\,\textrm{rk}(\tau)=\textrm{rk}(\sigma)-1,\sigma\subset\tau\}\textrm{ if rk}(\sigma)=r\\ &\emptyset\textrm{ otherwise}\end{aligned}\right.,$ (7)
and the up/down $r$-adjacencies $\mathcal{N}^{r}_{A,\uparrow}$ and $\mathcal{N}^{r}_{A,\downarrow}$ can be obtained analogously. So, returning to our previous example of a 2-dimensional complex, it is now straightforward to model a setting in which:
  • Employing only $\mathcal{N}^{1}_{I,\downarrow}$ (Fig. 4(iii)) allows edges to exchange messages with their bounding nodes but not triangles with their bounding edges.
  • Employing only $\mathcal{N}^{0}_{A,\uparrow}$ (Fig. 4(i)) allows nodes to exchange messages with their edge-connected nodes but not edges do not exchange messages with other edges that are part of their same faces.
@@
We use these notions to define a novel graph-based methodology for generating principled TDL architectures. Given a combinatorial complex $\mathcal{C}$ and a set $\mathcal{N_{C}}$ of neighborhoods, the method works as follows (see also Fig. 1):
  1. A.
    $\mathcal{C}$ is expanded into an ensemble of strictly augmented Hasse graphs—one for each $\mathcal{N}\in\mathcal{N}_{\mathcal{C}}$.
  2. B.
    Each strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}$ and the features of its cells are independently processed by a base model.
  3. C.
    An aggregation module $\bigotimes$ synchronizes the cell features across the different strictly augmented Hasse graphs (as the same cells can belong to multiple strictly augmented Hasse graphs).
This method enables an ensemble of synchronized models per layer— the $\omega_{\mathcal{N}}$s—each of them applied to a specific strictly augmented Hasse graph. [^3] Contrary to previous CCNN simulation works that leverage regular augmented graphs, in which one model is applied to a big graph.. Additionally, such a pipeline confers unprecedented flexibility in choosing a subset of neighborhoods of interest, allowing the consideration of per-rank neighborhoods within TDL. The rest of this section formalizes the architectures induced by this methodology and describes their theoretical properties.
@@
where $\mathbf{H}^{0}$ collects the initial features, and the update function $\phi$ is a learnable row-wise update function, i.e. $[\phi(\mathbf{A},\mathbf{B})]_{i}=\phi([\mathbf{A}]_{i},[\mathbf{B}]_{i})$. The neighborhood-dependent sub-module $\omega_{\mathcal{N}}:\mathbb{R}^{|\mathcal{C}_{\mathcal{N}}|\times F^{l}}\rightarrow\mathbb{R}^{|\mathcal{C}_{\mathcal{N}}|\times F^{l+1}}$, which we refer to as the neighborhood message function, is a learnable (matrix) function that takes as input the whole strictly augmented Hasse graph of the neighborhood, $\mathcal{G}_{\mathcal{N}}$ and the embeddings of the cells that are part of it, and gives as output a processed version of them. Finally, the inter-neighborhood aggregation module $\bigotimes$ synchronizes the possibly multiple neighborhood messages arriving on a single cell across multiple strictly augmented Hasse graphs into a single message. In this way, the embedding of a cell collects information about the whole relational structures induced by each (nonempty) neighborhood.
@@
Proposition 1.
Let $\mathcal{C}$ be a combinatorial complex. Let $\mathcal{N_{C}}$ be a collection of neighborhoods on $\mathcal{C}$. Then, there exists a GCCN that exactly reproduces the computation of a CCNN over $\mathcal{C}$ using $\mathcal{N_{C}}$. Proof. See Appendix A.1.
@@
Proposition 2.
A GCCN layer is cell permutation equivariant if the neighborhood message function is node permutation equivariant and the inter-neighborhood aggregator is cell permutation invariant. Proof. See Appendix A.2
3. Expressivity. The expressiveness of TDL models is tied to their ability to distinguish non-isomorphic graphs. Variants of the Weisfeiler-Leman (WL) test, like the cellular WL for cell complexes (Bodnar et al., 2021a), set upper bounds on their corresponding TDL models’ expressiveness, as the WL test does for GNNs (Xu et al., 2019a).
Proposition 3.
Let $\mathcal{C}$ be a combinatorial complex and $\mathcal{N_{C}}$ be a collection of neighborhoods. There exist GCCNs being as expressive as CCNNs. Proof. See Appendix A.3.
Given Proposition 1, GCCNs allow us to define general TDL models using any neighborhood message function $\omega_{\mathcal{N}}$, such as any GNN. Not only does this framework avoid having to approximate CCNN computations, as is the case in previous works  [^4] These models employ GNNs running on the whole augmented Hasse graph, i.e. a GCCN that, given a collection of neighborhoods $\mathcal{N_{C}}$, uses a single neighborhood $\mathcal{N}_{tot}$ defined, for a cell $\sigma$, as $\mathcal{N}_{tot}(\sigma)=\bigcup_{\mathcal{N}\in\mathcal{N_{C}}}\mathcal{N}(\sigma)$. (Jogl et al., 2022b; a; 2023), but it also enjoys the same permutation equivariance as regular CCNNs (Proposition 2). Differently from the work in (Hajij et al., 2023), the fact that GCCNs can have arbitrary neighborhood message functions implies that non message-passing TDL models can be readily defined (e.g., by using non message-passing models as neighborhood message functions). Moreover, the fact that the whole strictly augmented Hasse graphs are given as input enables also the usage of multi-layer GNNs as neighborhood message functions. To the best of our knowledge, GCCNs are the only objects in the literature that encompass all the above properties.
@@ 5 TopoTune
Our proposed methodology, together with its resulting GCCNs architectures, addresses the challenge of systematically generating principled, general TDL models. Here, we introduce TopoTune, a software module for defining and benchmarking GCCN architectures on the fly—a vehicle for accelerating and democratizing TDL research. TopoTune is made available as part of TopoBenchmark Telyatnikov et al. (2024) at github.com/geometric-intelligence/TopoBenchmark. This section details TopoTune’s main features.
@@
TopoTune introduces a new perspective on TDL through the concept of "neighborhoods of interest," enabling unprecedented flexibility in architectural design. Previously "fixed" components of CCNNs become hyperparameters of our framework. Even the choice of topological domain becomes a mere variable, representing a new paradigm in the design and implementation of TDL architectures.
@@
Using TopoTune, a practitioner can instantiate customized GCCNs simply by modifying a few lines of a configuration file. In fact, it is sufficient to specify $(i)$ a collection of per-rank neighborhoods $\mathcal{N}_{\mathcal{C}}$, $(ii)$ a neighborhood message function $\omega_{\mathcal{N}}$, and optionally $(iii)$ some architectural parameters—e.g. the number $l$ of GCCN layers. [^5] We provide a detailed pseudo-code for TopoTune module in Appendix B. For the neighborhood message function $\omega_{\mathcal{N}}$, the same configuration file enables direct import of models from standard PyTorch libraries, including PyTorch Geometric (Fey & Lenssen, 2019) and Deep Graph Library (Chen et al., 2020b). TopoTune’s simplicity provides both newcomers and TDL experts with an accessible tool for defining higher-order topological architectures.
@@
TopoTune is fully integrated into TopoBenchmark (Telyatnikov et al., 2024). This package provides standardized methods for handling combinatorial complex objects, including lifting procedures to lift them from graph-based data. TopoBenchmark also simplifies the training and benchmarking of GCCNs, offering a variety of ready-to-use tasks, datasets, and evaluation metrics. Together, TopoTune and TopoBenchmark provide all the necessary tools to define, train, test, and compare a wide range of novel GCCN architectures, bringing unprecedented versatility and standardization to accelerate TDL research.

### v3

@@ 3 Motivation and Related Works
The position paper (Veličković, 2022) proposed that any function over a higher-order domain can be computed via message passing over a transformed graph, but without specifying how to design GNNs that reproduce CCNNs. Hajij et al. (2023) showed that, given a combinatorial complex $\mathcal{C}$ and a collection of neighborhoods $\mathcal{N_{C}}$, a message-passing GNN that runs over the augmented Hasse graph $\mathcal{G}_{\mathcal{N_{C}}}$ is equivalent to a specific CCNN as in (4) running over $\mathcal{C}$ using: i) $\mathcal{N_{C}}$ as collection of neighborhoods; ii) same intra- and inter-aggregations, i.e., $\bigoplus$=$\bigotimes$ = ⨂; and iii) no rank- and neighborhood-dependent message functions, i.e., $\psi_{\mathcal{N},\textrm{rk}(\cdot)}=\psi\;\forall\mathcal{N}\in\mathcal{N_{C}}$.
Retaining expressivity, but not topological symmetry.
Jogl et al. (2022a, b) demonstrate that GNNs on augmented Hasse graphs $\mathcal{G}_{\mathcal{N_{C}}}$ are as expressive as CCNNs on $\mathcal{C}$ (using the WL criterion), suggesting that some CCNNs can be simulated with standard graph libraries.  [^2] The same authors generalize these ideas to non-standard message-passing GNNs (Jogl et al., 2024).. However, as the authors state, such GNNs do not structurally distinguish between cells of different ranks or neighborhoods, collapsing topological relationships. For instance, in a molecule (cellular complex), two bonds (edges) may simultaneously share multiple neighborhoods: lower-adjacent through a shared atom (node) and upper-adjacent through a shared ring (face). A GNN on $\mathcal{G}_{\mathcal{N_{C}}}$ collapses these distinctions, applying the same weights to all connections and losing the structural symmetries encoded in the domain. While this may suffice for preserving expressivity, it is inherently a very different computation than that of TDL models.
@@
Hypergraph neural networks have long relied on graph expansions (Telyatnikov et al., 2023), which has allowed the field to leverage advances in the graph domain and, by extension, a much wider breadth of models (Antelmi et al., 2023; Papillon et al., 2023). Most hypergraph models are expanded into graphs using the star (Zhou et al., 2006; Solé et al., 1996), the clique (Bolla, 1993; Rodríguez, 2002; Gibson et al., 2000), or the line expansion (Bandyopadhyay et al., 2020). As noted by Agarwal et al. (2006), many hypergraph learning algorithms leverage graph expansions.
The success story of hypergraph neural networks motivates further research on new graph-based expansions of CCNNs. These expansions could, at the same time, subsume current CCNNs and exploit progress in the GNN field. Therefore, returning to our core goal of accelerating and democratizing TDL while preserving its theoretical properties, we propose a two-part approach: a novel graph-based methodology able to generate general architectures (Section 4), and a lightweight software framework to easily and widely implement it (Section 5).
@@ 4 Generalized Combinatorial Complex Neural Networks
We propose Generalized Combinatorial Complex Neural Networks (GCCNs), a novel broad class of TDL architectures. GCCNs overcome the limitations of previous graph-based TDL architectures by leveraging the notions of strictly augmented Hasse graphs and per-rank neighborhoods.
Figure 3: Ensemble of strictly augmented Hasse Graphs. Given a complex $\mathcal{C}$ with neighborhood structure including both incidence and upper adjacency (left), this graph expansion (right) produces one augmented Hasse graph for each neighborhood.
@@
This graph expansion method (see Fig. 3) extends from the established definition of an augmented Hasse graph (see Fig. 2). Specifically, given a combinatorial complex $\mathcal{C}$ and a collection of neighborhood functions $\mathcal{N_{C}}$, we expand it into $|\mathcal{N_{C}}|$ graphs, each of them representing a neighborhood $\mathcal{N}\in\mathcal{N}_{\mathcal{C}}$. In particular, the strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}=(\mathcal{C}_{\mathcal{N}},\mathcal{E}_{\mathcal{N}})$ of a neighborhood $\mathcal{N}\in\mathcal{N}_{\mathcal{C}}$ is a directed graph whose nodes $\mathcal{C}_{\mathcal{N}}$ and edges $\mathcal{E}_{\mathcal{N}}$ are given by:
@@
Following the same arguments from (Hajij et al., 2023), a GNN over the strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}$ induced by $\mathcal{N}$ is equivalent to a CCNN running over $\mathcal{C}$ and using $\mathcal{N_{C}}=\{\mathcal{N}\}$ up to the (self-)update of the cells in $\mathcal{C}/\mathcal{C}_{\mathcal{N}}$.
@@
The standard definition of adjacencies and incidences given in Section 2 implies that they are applied to each cell regardless of its rank. For instance, consider a combinatorial complex of dimension two with nodes (0-cells), edges (1-cells), and faces (2-cells). Employing the down incidence $\mathcal{N}_{I,\downarrow}$ as in (1) means the edges must exchange messages with their endpoint nodes, and faces must exchange messages with the edges on their sides. It is impossible for edges to communicate while faces do not.
This limitation increases the computational cost of standard CCNNs while not always increasing the learning performance, as experiments will show. For this reason, we introduce per-rank neighborhoods, examples of which are depicted in Fig. 4. Formally, a per-rank neighborhood function $\mathcal{N}^{r}$ maps a cell $\sigma$ to the empty set if $\sigma$ is a cell of rank $r$. For example, the up/down $r$-incidences $\mathcal{N}^{r}_{I,\uparrow}$ and $\mathcal{N}^{r}_{I,\downarrow}$ are defined as
$\mathcal{N}^{r}_{I,\uparrow}(\sigma)=\begin{cases}\big{\{}\tau\in\mathcal{C}\,\big{|}\,\textrm{rk}(\tau)=\textrm{rk}(\sigma)+1,\,\sigma\subset\tau\big{\}},\\ \hskip 20.00003pt\text{if }\textrm{rk}(\sigma)=r,\\ \emptyset,\hskip 11.00008pt\text{otherwise},\end{cases}$
$\mathcal{N}^{r}_{I,\downarrow}(\sigma)=\begin{cases}\big{\{}\tau\in\mathcal{C}\,\big{|}\,\textrm{rk}(\tau)=\textrm{rk}(\sigma)-1,\,\sigma\subset\tau\big{\}},\\ \hskip 20.00003pt\text{if }\textrm{rk}(\sigma)=r,\\ \emptyset,\hskip 11.00008pt\text{otherwise}.\end{cases}$ (7)
and the up/down $r$-adjacencies $\mathcal{N}^{r}_{A,\uparrow}$ and $\mathcal{N}^{r}_{A,\downarrow}$ can be obtained analogously. So, it is now straightforward to model a setting in which employing only $\mathcal{N}^{1}_{I,\downarrow}$ (Fig. 4iii) allows edges to exchange messages with their bounding nodes but not triangles with their bounding edges.
@@
We use these notions to define a novel graph-based methodology for generating principled TDL architectures. Given a combinatorial complex $\mathcal{C}$ and a set $\mathcal{N_{C}}$ of neighborhoods, the method works as follows (see also Fig. 1): (i) $\mathcal{C}$ is expanded into an ensemble of strictly augmented Hasse graphs—one for each $\mathcal{N}\in\mathcal{N}_{\mathcal{C}}$. (ii) Each strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}$ and the features of its cells are independently processed by a base model. (iii) An aggregation module $\bigotimes$ synchronizes the cell features across the different strictly augmented Hasse graphs (as the same cells can belong to multiple strictly augmented Hasse graphs).
This method enables an ensemble of synchronized models per layer— the $\omega_{\mathcal{N}}$s—each of them applied to a specific strictly augmented Hasse graph. [^3] Contrary to past CCNN simulation works that apply a model to the singular, whole augmented Hasse graph.. The rest of this section formalizes the architectures induced by this methodology.
@@
where $\mathbf{H}^{0}$ collects the initial features, and the update function $\phi$ is a learnable row-wise update function, i.e., $[\phi(\mathbf{A},\mathbf{B})]_{i}=\phi([\mathbf{A}]_{i},[\mathbf{B}]_{i})$. The neighborhood-dependent sub-module $\omega_{\mathcal{N}}:\mathbb{R}^{|\mathcal{C}_{\mathcal{N}}|\times F^{l}}\rightarrow\mathbb{R}^{|\mathcal{C}_{\mathcal{N}}|\times F^{l+1}}$, which we refer to as the neighborhood message function, is a learnable (matrix) function that takes as input the whole strictly augmented Hasse graph of the neighborhood, $\mathcal{G}_{\mathcal{N}}$ and the embeddings of the cells that are part of it, and gives as output a processed version of them. Finally, the inter-neighborhood aggregation module $\bigotimes$ synchronizes the possibly multiple neighborhood messages arriving on a single cell across multiple strictly augmented Hasse graphs into a single message. GCCNs enjoy increased flexibility over CCNS (eq. 4) as their neighborhoods are allowed to be rank-dependent and their $\omega_{\mathcal{N}}$’s are not necessarily message-passing.
@@
Proposition 4.1.
Let $\mathcal{C}$ be a combinatorial complex. Let $\mathcal{N_{C}}$ be a collection of neighborhoods on $\mathcal{C}$. Then, there exists a GCCN that exactly reproduces the computation of a CCNN over $\mathcal{C}$ using $\mathcal{N_{C}}$.
@@
Proposition 4.2.
A GCCN layer is cell permutation equivariant if the neighborhood message function is node permutation equivariant and the inter-neighborhood aggregator is cell permutation invariant.
3. Expressivity. The expressiveness of TDL models is tied to their ability to distinguish non-isomorphic graphs. Variants of the Weisfeiler-Leman (WL) test, like the cellular WL for cell complexes (Bodnar et al., 2021a), set upper bounds on their corresponding TDL models’ expressiveness, as the WL test does for GNNs (Xu et al., 2019b).
Proposition 4.3.
GCCNs are strictly more expressive than CCNNs.
Proofs are provided in Appendix B.1, B.2, and B.3, respectively.
Given Proposition 4.1, GCCNs allow us to define general TDL models using any neighborhood message function $\omega_{\mathcal{N}}$, such as any GNN. Not only does this framework avoid having to approximate CCNN computations, as is the case in previous works  [^4] These models employ GNNs running on one augmented Hasse graph, i.e. a GCCN that, given a collection of neighborhoods $\mathcal{N_{C}}$, uses a single neighborhood $\mathcal{N}_{tot}$ defined, for a cell $\sigma$, as $\mathcal{N}_{tot}(\sigma)=\bigcup_{\mathcal{N}\in\mathcal{N_{C}}}\mathcal{N}(\sigma)$. (Jogl et al., 2022b, a, 2023), but it also enjoys the same permutation equivariance as regular CCNNs (Proposition 4.2). We show in Appendix C that the resulting time complexity of a GCCN is a compromise between a typical GNN and a CCNN. Differently from the work in (Hajij et al., 2023), the fact that GCCNs can have arbitrary neighborhood message functions implies that non message-passing TDL models can be readily defined. To the best of our knowledge, GCCNs are the only objects in the literature that encompass all the above properties.
@@ 5 TopoTune
Our proposed methodology, together with its resulting GCCNs architectures, addresses the challenge of systematically generating principled, general TDL models. Here, we introduce TopoTune, a software module for defining and benchmarking GCCN architectures on the fly—a vehicle for accelerating TDL research. All code and experiment scripts are provided at github.com/geometric-intelligence/TopoBenchmark.
@@
TopoTune introduces a new perspective on TDL through the concept of “neighborhoods of interest,” enabling unprecedented flexibility in architectural design. Previously fixed components of CCNNs, such as choice of topological domain, become hyperparameters of our framework.
@@
Using TopoTune, a practitioner can instantiate customized GCCNs simply by modifying a few lines of a configuration file. In fact, it is sufficient to specify $(i)$ a collection of per-rank neighborhoods $\mathcal{N}_{\mathcal{C}}$, $(ii)$ a neighborhood message function $\omega_{\mathcal{N}}$, and optionally $(iii)$ some architectural parameters—e.g., the number $l$ of GCCN layers. [^5] We provide a detailed pseudo-code for TopoTune module in Appendix D. For the neighborhood message function $\omega_{\mathcal{N}}$, the same configuration file enables direct import of models from PyTorch libraries such as PyTorch Geometric (Fey & Lenssen, 2019) and Deep Graph Library (Chen et al., 2020b). TopoTune’s simplicity provides both newcomers and experts with an accessible tool for defining topological architectures.
@@
TopoTune is fully integrated into TopoBenchmark (Telyatnikov et al., 2024), a comprehensive package offering a wide range of standardized methods and tools for TDL. Practitioners can access ready-to-use models, training pipelines, tasks, and evaluation metrics, including leading open-source models from TopoX (Hajij et al., 2024b). In addition, TopoBenchmark features the largest collection of topological liftings currently available—transformations that map graph datasets into higher-order topological domains. Together, TopoBenchmark and TopoTune organize the vast design space of TDL into an accessible framework, providing unparalleled versatility and standardization for practitionners.